OpenStudy (anonymous):

simplify the following radical please show work 1. √k^5/5√18k

OpenStudy (anonymous):

\[\frac{ \sqrt(k^5)}{ 5\sqrt{18k} }\] Is this right?

OpenStudy (jdoe0001):

\(\bf \cfrac{\sqrt{k^5}}{5\sqrt{18k}}\implies \cfrac{\sqrt{k^2\cdot k^2\cdot k}}{5\sqrt{3^2\cdot 2\cdot k}}\implies \cfrac{\sqrt{k^2}\cdot \sqrt{k^2}\cdot \sqrt{k}}{5\sqrt{3^2}\cdot \sqrt{2}\cdot \sqrt{k}}\)

OpenStudy (anonymous):

what is the answer?